A quadratic equation has roots 1+-2i.What is the equation?
Answer by ankor@dixie-net.com(22740)

A quadratic equation has roots 1+-2i.What is the equation?
:
Sort of like completing the square in reverse:
:
x = 1 +/- 2i
:
x - 1 = +/- 2i; subtract 1 from both sides
:
(x-1)^2 = (+/-2i)^2; square both sides
:
x^2 - 2x + 1 = 4(i^2); FOIL the left side
:
x^2 - 2x + 1 = 4(-1); i^2 = -1
:
x^2 - 2x + 1 = -4
:
x^2 - 2x + 1 + 4 = 0; add 4 to both sides
:
x^2 - 2x + 5 = 0; is the quadratic equation
:
You can check this by completing the square, and see if you get the given roots